the age of the older of 2 brothers is double of that of the younger . however 6 years ago it was 4 times that of the younger. what are their ages in years?
Answer by mananth(16946)


the age of the older of 2 brothers is double of that of the younger . however 6 years ago it was 4 times that of the younger.
let younger brother's age be x
Elder brother's age is 2x
6 years ago
younger's age = x-6
Elder's age was 2x-6
Relation
2x-6 = 4(x-6)
2x-6 = 4x-24
2x=18
x=9
2x =18